# On the mixed monotonicity of polynomial functions

### v1

@@ V Examples
In this section four examples are provided. In all of the examples, Convex.jl [13] with the SCS solver [14] was used to solve the semidefinite program (8).
V-A Example 1: Comparison to decomposition function based on Jacobian bounds
@@
To illustrate the decomposition function, consider Fig. 1 where for $z=-5$ to $z=5$, $g_{1}(z+1,z-0.5)$ is plotted in red and $g_{1}(z-0.5,z+1)$ is plotted in blue. Additionally, $p_{1}(z+1),p_{1}(z+0.5),p_{1}(z),p_{1}(z-0.25),$ and $p_{1}(z-0.5)$ are plotted in gray. Since $g_{1}$ is a decomposition function, then $g_{1}(z-0.5,z+1)\leq p_{1}(z+1),p_{1}(z+0.5),p_{1}(z),p_{1}(z-0.25),p_{1}(z-0.5)\leq g_{1}(z+1,z-0.5)$ for all $z\in\mathbb{R}$. This can clearly be seen in Fig. 1 by the fact that all of the gray lines lie in between the red and blue lines. It appears that the decomposition function is less tight as $\|z\|$ increases.
@@
In Fig. 2, for $z=-1$ to $z=1$, $g_{2}(z+1,z-0.5)$ is plotted in green and $g_{2}(z-0.5,z+1)$ is plotted in magenta along with $g_{1}(z+1,z-0.5)$ in red and $g_{1}(z-0.5,z+1)$ in blue. The gray lines are the same as shown in Fig. 1. In addition to being a global decomposition function, it is clear from Fig. 2 that the polynomial decomposition function is much tighter than the decomposition function derived from Jacobian bounds.
V-B Example 2: Tightness of polynomial decomposition functions derived from different objective functions
@@
Fig. 3 shows a comparison of the decomposition function $g_{3}$ given in (13) and $g_{4}$ given in (15) by plotting $g_{3}(z+0.25,z-0.3)$, $g_{3}(z-0.3,z+0.25)$, $g_{4}(z+0.25,z-0.3)$ and $g_{4}(z-0.3,z+0.25)$ for $z=-1.5$ to $z=1.5$ in red, blue, green, and magenta, respectively. It can be clearly seen that the interval $[g_{3}(z-0.3,z+0.25),g_{3}(z+0.25,z-0.3)]$ is tighter than the interval $[g_{4}(z-0.3,z+0.25),g_{4}(z+0.25,z-0.3)]$ for all $z=-1.5$ to $z=1.5$.
V-C Example 3: Checking monotonicity of a polynomial
@@
V-D Example 4: Reachable set computation
Consider the following discrete-time system inspired by the example in [15, §VI.B]:
@@ V-D Example 4: Reachable set computation
The exact reachable set after $N$ steps is defined as follows (cf. [15, Definition 4]):
[…]
@@
In Fig. 5, the over-approximation of the reachable set computed by propagating (V-D) for the system (V-D) is shown where $\underline{\xi}_{0}=x[0]=\overline{\xi}_{0}=0$ in (19), i.e. reachability from the origin. $\overline{x}[k]$ is shown in red and $\underline{x}[k]$ is shown in blue. Sample trajectories of (V-D) with randomly generated inputs that satisfy the bounds (18) are shown in gray. All of the gray lines are contained within the red and blue lines, i.e. for each trajectory, $x[k]$ belongs the interval $\left[\underline{x}[k],\overline{x}[k]\right]$ for all $k=0,1,2,\dots$. Moreover, it is clear from Fig. 5 that the over-approximation is a tight over-approximation in the sense that the interval $\left[\underline{x}[k],\overline{x}[k]\right]$ is almost the smallest interval that contains all of the gray trajectories.
Figure 5: Depiction of the over-approximation of the reachable set from the origin computed by propagating (V-D). The gray lines are sample trajectories of (V-D) with randomly generated inputs that satisfy the bounds (18).
VI Conclusions
@@
There are a few avenues to pursue future work. Firstly, an important application would be to consider computing decomposition functions for polynomials with uncertain coefficients. This would be invaluable for constructing interval observers for uncertain polynomial systems. Secondly, further investigation into how to compute the tightest decomposition function should be pursued. Lastly, the results of this paper only pertain to univariate polynomials, so future work should be dedicated to constructing polynomial decompositions for multivariate polynomials or to showing the conditions under which multivariate polynomials have polynomial decomposition functions.

### v2

@@ 6 Examples
In this section five examples are provided. In the examples, Convex.jl [^2] https://jump.dev/Convex.jl/stable/ with the SCS solver was used to solve the semidefinite program (19). To verify nonnegativity of polynomials using sum of squares techniques, the package SumOfSquares.jl [^3] https://jump.dev/SumOfSquares.jl/stable/ was used.
6.1 Example 1: Comparison to decomposition function based on Jacobian bounds
@@
To illustrate the decomposition function, consider Fig. 1 where for z=5z=-5 to $z=5$, $g_{1}(z+1,z-0.5)$ is plotted in red and $g_{1}(z-0.5,z+1)$ is plotted in blue. Additionally, $p_{1}(z+1),p_{1}(z+0.5),p_{1}(z),p_{1}(z-0.25),$ and $p_{1}(z-0.5)$ are plotted in gray. Since $g_{1}$ is a decomposition function, then $g_{1}(z-0.5,z+1)\leq p_{1}(z+1),p_{1}(z+0.5),p_{1}(z),p_{1}(z-0.25),p_{1}(z-0.5)\leq g_{1}(z+1,z-0.5)$ for all $z\in\mathbb{R}$. This can clearly be seen in Fig. 1 by the fact that all of the gray lines lie in between the red and blue lines. It appears that the decomposition function is less tight as $\|z\|$ increases.
@@
In Fig. 2, for z=1z=-1 to $z=1$, $g_{2}(z+1,z-0.5)$ is plotted in green and $g_{2}(z-0.5,z+1)$ is plotted in magenta along with $g_{1}(z+1,z-0.5)$ in red and $g_{1}(z-0.5,z+1)$ in blue. The gray lines are the same as shown in Fig. 1. In addition to being a global decomposition function, it is clear from Fig. 2 that the polynomial decomposition function is much tighter than the decomposition function derived from Jacobian bounds.
6.2 Example 2: Tightness of polynomial decomposition functions derived from different objective functions
@@
Fig. 3 shows a comparison of the decomposition function $g_{3}$ given in (26) and $g_{4}$ given in (28) by plotting $g_{3}(z+0.25,z-0.3)$, $g_{3}(z-0.3,z+0.25)$, $g_{4}(z+0.25,z-0.3)$ and $g_{4}(z-0.3,z+0.25)$ for z=1.5z=-1.5 to $z=1.5$ in red, blue, green, and magenta, respectively. It can be clearly seen that the interval $g_{3}(z-0.3,z+0.25),g_{3}(z+0.25,z-0.3)]$ is tighter than the interval $g_{4}(z-0.3,z+0.25),g_{4}(z+0.25,z-0.3)]$ for all z=1.5z=-1.5 to $z=1.5$.
6.3 Example 3: Checking monotonicity of a polynomial
@@
6.4 Example 4: Multivariate polynomial function
Consider the following multivariate monomial:
$p_{m}(x)=x_{1}x_{2}.$ (30)
According to Theorem 2, $p_{m}$ is mixed monotone globally over $\mathbb{R}^{2}$ and has a polynomial decomposition function. The monomial $p_{m}$ is the product of two mixed monotone functions ($p_{m}(x)=f_{1}(x)f_{2}(x)$ where $f_{1}(x)=e_{1}^{\top}x=x_{1}$ and $f_{2}(x)=e_{2}^{\top}x=x_{2}$). Therefore, using Lemma 4, $p_{m}$ has the following mixed monotone decomposition function:
$g_{m}(x,y)=-0.3536y_{2}-0.3536y_{1}+0.3536x_{2}+0.3536x_{1}$
$+0.25y_{1}y_{2}+0.25x_{2}y_{1}+0.25x_{1}y_{2}+0.25x_{1}x_{2}$
$-0.05893y_{2}^{3}-0.08839y_{1}y_{2}^{2}-0.08839y_{1}^{2}y_{2}-0.05893y_{1}^{3}$
$+0.08839x_{2}y_{1}^{2}-0.08839x_{2}^{2}y_{1}+0.05893x_{2}^{3}$
$+0.08839x_{1}y_{2}^{2}+0.08839x_{1}x_{2}^{2}-0.08839x_{1}^{2}y_{2}$
$+0.08839x_{1}^{2}x_{2}+0.058934x_{1}^{3}.$ (31)
To compute this decomposition function, $g_{sq}(x,y)$ was chosen similarly to Example 1 (i.e. $g_{sq}(x,y)=g_{1}(x,y)-1$ where $g_{1}$ is given in (24)). Now it is shown that $g_{m}$ satisfies all three properties of a decomposition function for $p_{m}$.
  1. 1.
    The first property can be verified by substitution to find $g_{m}(x,x)=p_{m}(x)$.
  2. 2.
    Now it will be shown that $g_{m}$ is increasing in its first argument by showing that the partial derivative with respect to $x_{1}$ and $x_{2}$ can be written as sums of squares:
    1. (a)
      The partial derivative $\frac{\partial g_{m}}{\partial x_{1}}(x,y)=0.3536+0.25y_{2}+0.25x_{2}+0.08839y_{2}^{2}+0.08839x_{2}^{2}-0.1768x_{1}y_{2}+0.1768x_{1}x_{2}+0.1768x_{1}^{2}=(-0.5946-0.2102y_{2}-0.2102x_{2})^{2}+(-0.2102y_{2}-0.2102x_{2}-0.4204x_{1})^{2}\geq 0$ for all $(x,y)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$.
    2. (b)
      The partial derivative $\frac{\partial g_{m}}{\partial x_{2}}(x,y)=(-0.5946-0.2102y_{1}-0.2102x_{1})^{2}+(0.2102y_{1}-0.4204x_{2}-0.2102x_{1})^{2}\geq 0$ for all $(x,y)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$.
  3. 3.
    Finally, $g_{m}$ is shown to be decreasing in its second argument by showing that the partial derivatives with respect to $y_{1}$ and $y_{2}$ can be expressed as the negations of sums of squares:
    1. (a)
      The partial derivative $\frac{\partial g_{m}}{\partial y_{1}}(x,y)=-(-0.5946+0.2102y_{2}+0.2102x_{2})^{2}-(0.2102y_{2}+0.4204y_{1}-0.2102x_{1})^{2}_{2}\leq 0$ for all $(x,y)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$.
    2. (b)
      The partial derivative $\frac{\partial g_{m}}{\partial y_{2}}(x,y)=-(-0.5946+0.2102y_{1}+0.2102x_{1})^{2}-(0.4204y_{2}-0.2102y_{2}+0.2102_{1}x_{1})^{2}\leq 0$ for all $(x,y)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$.
6.5 Example 5: Reachable set computation
Consider the following discrete-time system inspired by the example in [1, §VI.B]:
@@ 6.5 Example 5: Reachable set computation
The exact reachable set after $N$ steps is defined as follows (cf. [1, Definition 4]):
$\mathcal{R}_{N}=\{$ $x[N]\in\mathbb{R}:x[k+1]=f(x[k])+u[k],$
@@
In Fig. 5, the over-approximation of the reachable set computed by propagating (35) for the system (32) is shown where $\underline{\xi}_{0}=x[0]=\overline{\xi}_{0}=0$ in (34), i.e. reachability from the origin. $\overline{x}[k]$ is shown in red and $\underline{x}[k]$ is shown in blue. Sample trajectories of (32) with randomly generated inputs that satisfy the bounds (33) are shown in gray. All of the gray lines are contained within the red and blue lines, i.e. for each trajectory, $x[k]$ belongs the interval $\left[\underline{x}[k],\overline{x}[k]\right]$ for all $k=0,1,2,\dots$. Moreover, it is clear from Fig. 5 that the over-approximation is a tight over-approximation in the sense that the interval $\left[\underline{x}[k],\overline{x}[k]\right]$ is almost the smallest interval that contains all of the gray trajectories.
Figure 5: Depiction of the over-approximation of the reachable set from the origin computed by propagating (35). The gray lines are sample trajectories of (32) with randomly generated inputs that satisfy the bounds (33).
7 Conclusions
@@
There are a few avenues to pursue future work. Firstly, an important application would be to consider computing decomposition functions for polynomials with uncertain coefficients. This would be invaluable for constructing interval observers for uncertain polynomial systems. Secondly, further investigation into how to compute the tightest decomposition function should be pursued. More work should be dedicated to constructing decomposition functions for multivariate polynomials.